On January 1, 2016, Oriole Company issued 10-year, $96,000 face value, 6% bonds at par (interest payable annually on January 1). Each $900 bond is convertible into 32 shares of Oriole $2 par value common stock. The company has had 9,000 shares of common stock (and no preferred stock) outstanding throughout its life. None of the bonds have been converted as of the end of 2017. Oriole also has adopted a stock-option plan that granted options to key executives to purchase 5,400 shares of the company’s common stock. The options were granted on January 2, 2016, and were exercisable 2 years after the date of grant if the grantee was still an employee of the company (the service period is 2 years). The options expired 6 years from the date of grant. The option price was set at $5, and the fair value option-pricing model determines the total compensation expense to be $21,000. All of the options were exercised during the year 2018: 2,700 on January 3 when the market price was $7, and 900 on May 1 when the market price was $8 a share. (Ignore all tax effects.)

Prepare the journal entry Oriole would have made on January 1, 2016, to record the issuance of the bonds.

Prepare the journal entry to record interest expense and compensation expense in 2017.

Oriole’s net income in 2017 was $29,700. Compute basic and diluted earnings per share for Oriole for 2017. Oriole’s average stock price was $6 in 2017.

Assume that 75 percent of the holders of Oriole’s convertible bonds convert their bonds to stock on June 30, 2018,when Oriole’s stock is trading at $8 per share. Oriole pays $2 per bond to induce bondholders to convert. Preparethe journal entry to record the conversion.
Answer #1
a. Journal entry for issuance of Bonds:
Date Particulars Dr/Cr Amount Amount
1-Jan-16 Bank A/c Dr $     96,000
To 6% Convertible bonds A/c Cr $ 96,000
(Being 6% Convertible bonds issued at par)
b. Journal entry for interest expense and compensation expense:
Date Particulars Dr/Cr Amount Amount
1-Jan-17 Interest A/c Dr $       5,760
To Bank A/c Cr $ 5,760
(Being Interest expense accounted)
2-Jan-17 Employee Compensation expense A/c Dr $     21,000
To Employee Stock Option A/c Cr $ 21,000
(Being employee compensation expense accounted)
c. Basic Earnings per share
= Net Income for the year/No of shares at the end
= 29700/9000
= $         3.30
Diluted Earnings per share
= (Net Income)/ (Weighted average no of shares at the end + Other convertible instruments)
= (29700)/((9000/12*12)+($96000/$900*32))
= $         2.39
d. Conversion of bonds
Percentage of bonds converted 75%
Value of bonds converted = $96000*75% = $72,000
No of bonds converted = $72000/$900 = 80bonds
Each bond is converted into 32 shares
Therefore, no of shares converted = 32*80 shares= 2560shares
Value of shares converted = 2560shares* $2each = $5,120